OpenStudy (anonymous):

Evaluate. 3*1.5-2.5 over 4(8.62-6.37)

OpenStudy (wendy.ivette11714):

Do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (wendy.ivette11714):

Is this your question?:\[\frac{ (3)(1.5)−2.5 }{ 4(8.62−6.37) }\]

OpenStudy (anonymous):

yes

OpenStudy (wendy.ivette11714):

K

OpenStudy (wendy.ivette11714):

\[\frac{ (3)(1.5)−2.5 }{ 4(8.62−6.37) }\] \[= \frac{ 4.5−2.5 }{ 4(8.62−6.37) }\] \[= \frac{ 2 }{ 4(8.62−6.37) }\] \[= \frac{ 2 }{ (4)(2.25) }\] Answer: \[\frac{ 2 }{ 9 }~or~0.222222\]

OpenStudy (anonymous):

thank u do u want a medal

OpenStudy (wendy.ivette11714):

Your welcome. You already gave me one :)